# A martingale approach to noncommutative stochastic calculus

### v2

@@ 4.2 Integration against $L^{2}$-decomposable processes
    $\int_{0}^{\boldsymbol{\cdot}}K[\mathrm{d}Y]=\int_{0}^{\boldsymbol{\cdot}}KH[\mathrm{d}X],$
@@ 5.1 Itô product rule and reduction to martingales
$\mathbb{L}^{2}\text{-}\lim_{\Pi\in\mathcal{P}_{\mathbb{R}_{+}}}\mathrm{RS}_{\Pi}^{X,Y}(\Lambda)$ $=\Lambda[X,Y]-\Lambda(0)[X(0),Y(0)]-\int_{0}^{\boldsymbol{\cdot}}\mathrm{d}\Lambda(t)[X(t),Y(t)]$
$\hskip 39.83385pt-\int_{0}^{\boldsymbol{\cdot}}\Lambda(t)[\mathrm{d}X(t),Y(t)]-\int_{0}^{\boldsymbol{\cdot}}\Lambda(t)[X(t),\mathrm{d}Y(t)].$ (5.1)
@@
$\mathbb{L}^{r}\text{-}\lim_{(\Pi,\xi)\in\mathcal{P}_{\mathbb{R}_{+}}^{*}}\sum_{t\in\Pi}\Lambda(t_{*})[X(t\wedge\cdot)-X(t_{-}\wedge\cdot),Y(t\wedge\cdot)-Y(t_{-}\wedge\cdot)]=0,$
where $t_{*}=\xi(t)$ per Notation 4.1. In particular, $\mathbb{L}^{r}\text{-}\lim_{\Pi\in\mathcal{P}_{\mathbb{R}_{+}}}\mathrm{RS}_{\Pi}^{X,Y}(\Lambda)=0$.
  • Proof.
    Let $t\geq 0$, and define $C_{t}\coloneqq\sup\{\lVert\Lambda(s)\rVert_{p,q;r}:0\leq s\leq t\}<\infty$. If $(\Pi,\xi)\in\mathcal{P}_{\mathbb{R}_{+}}^{*}$, then
@@ 5.2 Construction of quadratic covariation
    $\Bigg{\|}\sum_{s\in\pi}\Lambda(s_{*})[\Delta_{s}M,\Delta_{s}N]\Bigg{\|}_{1}\leq\int_{0}^{t}\lVert\Lambda\rVert_{2,2;1}^{(\pi,\xi)}\,\mathrm{d}\kappa_{M,N}$ for all $(\pi,\xi)\in\mathcal{P}_{[0,t]}^{*}$.
@@
    $=\sum_{s\in\pi}\lVert\Lambda(s_{*})\rVert_{2,2;1}\kappa_{M,N}((s_{-},s])=\int_{0}^{t}\lVert\Lambda\rVert_{2,2;1}^{(\pi,\xi)}\,\mathrm{d}\kappa_{M,N},$
@@
    Now, if $|\pi|\to 0$, then $s_{+}\searrow s$. Thus, by the continuity of $X$ and $Y$, $\Lambda(s)[\Delta_{s_{+}}X,\Delta_{s_{+}}Y]\to 0$ in $L^{1}(\mathbb{E}_{{\mathcal{C}}})$ as $|\pi|\to 0$. It therefore follows, by taking $|\pi|\to 0$, from Theorem 5.7 and Definition 5.8 that
@@ 5.3 Tools to calculate quadratic covariations
    For the third claim, note that if $\Lambda$ is $\lVert\cdot\rVert_{2,2;1}$-continuous, $t\geq 0$, and $(\Pi,\xi)\in\mathcal{P}_{\mathbb{R}_{+}}^{*}$, then
    $\sup_{0\leq s\leq t}\delta_{(\Pi,\xi)}(s)$ $\coloneqq\sup_{0\leq s\leq t}\Bigg{\|}\sum_{r\in\Pi}\Lambda(r_{*})[\Delta_{r}M^{s},\Delta_{r}N^{s}]-\mathrm{RS}_{\Pi}^{M,N}(\Lambda)(s)\Bigg{\|}_{1}$
@@
    $=\kappa_{M,N}((0,t])\sup_{u,v\leq t:|u-v|\leq|\Pi|}\lVert\Lambda(u)-\Lambda(v)\rVert_{2,2;1}\xrightarrow[(\Pi,\xi)\in\mathcal{P}_{\mathbb{R}_{+}}^{*}]{|\Pi|\to 0}0$
@@ 5.4 Application: Burkholder–Davis–Gundy inequalities
Moreover, we may take $\alpha_{2}=\beta_{2}=1$, and $\lVert x\rVert_{\mathcal{H}^{2}(\mathcal{M})}^{2}=\tau\big{[}x_{0}^{*}x_{0}+\sum_{n=1}^{N}(x_{n}-x_{n-1})^{*}(x_{n}-x_{n-1})\big{]}$.
@@
    (ii) First, note that $a\in L^{q}(\mathbb{E})$ and $\lVert a\rVert_{q}\leq\liminf_{n\to\infty}\lVert a_{n}\rVert_{q}\leq\sup\{\lVert a_{n}\rVert_{q}:n\in\mathbb{N}\}<\infty$ by the first item. For the second claim, the case $p=1$ is obvious, so we assume $p>1$. If $b\in\mathcal{A}$ and $M>0$, then
@@
Moreover, if $p=2$, then $\lVert M(t)\rVert_{2}^{2}=\mathbb{E}\big{[}[M^{*},M]_{t}^{P}\big{]}=\mathbb{E}\big{[}[M,M^{*}]_{t}^{P}\big{]}$.
  • Proof.
@@ 5.5 Examples
[…]
    $=\mu_{M,N}((s,t])\,\mathbb{E}[\Xi[M(v)-M(u),N(v)-N(u)]\mid\mathcal{A}_{u}]\,\text{ and}$ (5.15)
@@
    $\mathbb{E}[\Xi[M(t)-M(s)$ $,N(t)-N(s)]\mid\mathcal{A}_{s}]$
    $=\frac{\mathbb{E}[\Xi[M(s+r)-M(s),N(s+r)-N(s)]\mid\mathcal{A}_{s}]}{\mu_{M,N}((s,s+r])}\,\mu_{M,N}((s,t])\quad(s<t,\;\Xi\in\mathcal{T}_{2,s}).$ (5.17)
@@
Above, we have written $X(t)\coloneqq X(t,\cdot)$ and considered $\mathbb{R}_{+}\ni t\mapsto X(t)\in L^{2}(\tau_{n})$ as a noncommutative $L^{2}$-martingale in the usual way. Now, by the “magic formulas” ([28, §3.1]),
@@ 6.1 Adapted $C^{k}$ maps
    if $i=0,\ldots,k$, $t\geq 0$, and $(a,b)\in(\mathcal{A}_{t}\times\mathcal{B}_{t})\cap U$, then $D_{1}^{i}F(a,b)\in\mathcal{F}_{i,t}(\mathbb{E}_{{\mathcal{A}}},\ldots,\mathbb{E}_{{\mathcal{A}}};\mathbb{E}_{{\mathcal{C}}})$; and
@@
    if $j=0,\ldots,\ell$, $t\geq 0$, and $(a,b)\in(\mathcal{A}_{t}\times\mathcal{B}_{t})\cap U$, then $D_{2}^{j}F(a,b)\in\mathcal{F}_{j,t}(\mathbb{E}_{{\mathcal{B}}},\ldots,\mathbb{E}_{{\mathcal{B}}};\mathbb{E}_{{\mathcal{C}}})$.
@@
where $S_{k}$ is the symmetric group on $k$ letters. Therefore, by Proposition 3.9, $F\in C_{a}^{\infty}(\mathcal{U};\mathcal{A})$.
@@
$\big{(}\partial_{\mathbf{b}_{k}}\cdots\partial_{\mathbf{b}_{1}}f_{*}(\mathbf{a}))(\omega)=\partial_{\mathbf{b}_{k}(\omega)}\cdots\partial_{\mathbf{b}_{1}(\omega)}f(\mathbf{a}(\omega))\qquad\big{(}\mathbf{a},\mathbf{b}_{1},\ldots,\mathbf{b}_{k}\in\mathcal{A}_{n,\beta}^{d},\text{ a.e.\ }\omega\in\Omega\big{)}.$ (6.2)
@@ 6.3 Examples: Trace smooth maps
The closure of $\{P_{{(\mathcal{A},\mathbb{E})}}:P\in(\mathrm{TrP}_{n,k,d}^{*})^{m}\}$ in the space $BC_{\operatorname{loc}}(\mathcal{A}_{\beta}^{n};\mathbb{B}_{k}(\mathcal{A}_{\gamma}^{d};\mathcal{A}^{m}))$ is contained in $C_{\mathbb{E}}(\mathcal{A}_{\beta}^{n};\mathbb{B}_{k}(\mathcal{A}_{\gamma}^{d};\mathcal{A}^{m}))$. In particular, if $F\in C^{k}(\mathcal{A}_{\beta}^{n};\mathcal{A}^{m})$ is such that for all $i=0,\ldots,k$, the $i^{\text{th}}$ derivative $D^{i}F$ belongs to the closure of $\{P_{{(\mathcal{A},\mathbb{E})}}:P\in(\mathrm{TrP}_{n,i,(n,\ldots,n)}^{*})^{m}\}$ in the space $BC_{\operatorname{loc}}(\mathcal{A}_{\beta}^{n};\mathbb{B}_{i}((\mathcal{A}_{\beta}^{n})^{i};\mathcal{A}^{m}))$, then $F\in C_{\mathbb{E}}^{k}(\mathcal{A}_{\beta}^{n};\mathcal{A}^{m})$. Consequently, the tracial noncommutative $C^{k}$ functions introduced and studied in [37] provide examples of elements of $C_{\mathbb{E}}^{k}(\mathcal{A}_{\operatorname{sa}}^{n};\mathcal{A}^{m})$. In the next section, we give examples of this kind that arise from the functional calculus (i.e., operator functions); see Remark 6.40.
@@ 6.4 Examples: Scalar functions
If, in addition, $m=k+1$, $\mathbf{a}=(a_{1},\ldots,a_{k+1})\in\mathcal{M}_{\operatorname{sa}}^{k+1}$, and $\Omega_{j}=\sigma(a_{j})$ ($j=1,\ldots,k+1$), then
@@
By an elementary induction argument, $f^{[k]}(\lambda_{1},\ldots,\lambda_{k+1})=\sum_{i=1}^{k+1}f(\lambda_{i})\prod_{j\neq i}(\lambda_{i}-\lambda_{j})^{-1}$ for all distinct $\lambda_{1},\ldots,\lambda_{k+1}\in S$. In particular, $f^{[k]}$ is symmetric in its arguments. Now, we state a useful expression for $f^{[k]}$ when $f\in C^{k}(\mathbb{R})$ or when $f\colon\mathbb{C}\to\mathbb{C}$ is entire; see [49, Prop. 2.1.3(ii)] for a proof.
@@ Appendix A $L^{p}$ spaces of $\mathrm{C}^{*}$-probability spaces
Let $p\in[1,\infty)$, and write $\lVert a\rVert_{p}\coloneqq\mathbb{E}[|a|^{p}]^{1/p}$ and $\lVert a\rVert_{\infty}\coloneqq\lVert a\rVert$ for all $a\in\mathcal{A}$.
@@
    If $a\in\mathcal{A}$, then $\lVert a\rVert_{p}=\sup\{|\mathbb{E}[ab]|:b\in\mathcal{A},\,\lVert a\rVert_{q}\leq 1\}$. If, in addition, $1<p,q<\infty$, then the map $\mathcal{A}\ni a\mapsto(b\mapsto\tilde{\mathbb{E}}[ab])\in L^{p}(\mathcal{A},\mathbb{E})^{*}$ extends to an isometric isomorphism $L^{q}(\mathcal{A},\mathbb{E})\to L^{p}(\mathcal{A},\mathbb{E})^{*}$.
@@ Appendix B Notation index
$\ell_{(\operatorname{loc})}^{\infty}(\mathbb{R}_{+};\mathcal{V})$ (locally) bounded functions $\mathbb{R}_{+}\to\mathcal{V}$; pp. 4.1 & 5.7
@@ Acknowledgments.
D. A. Jekel acknowledges support from NSF grant DMS-2002826. T. A. Kemp acknowledges support from NSF grants DMS-2055340 and DMS-1800733. E. A. Nikitopoulos acknowledges support from NSF grant DGE-2038238 and partial support from NSF grant DMS-2055340.

### v3

@@ 4.2 Integration against $L^{2}$-decomposable processes
    $\int_{0}^{\boldsymbol{\cdot}}K[\mathrm{d}Y]=\int_{0}^{\boldsymbol{\cdot}}KH[\mathrm{d}X]$
@@ 5.1 Itô product rule and reduction to martingales
$\begin{split}\mathbb{L}^{2}\text{-}\lim_{\Pi\in\mathcal{P}_{\mathbb{R}_{+}}}\mathrm{RS}_{\Pi}^{X,Y}(\Lambda)=\Lambda[X&,Y]-\Lambda(0)[X(0),Y(0)]-\int_{0}^{\boldsymbol{\cdot}}\mathrm{d}\Lambda(t)[X(t),Y(t)]\\ &-\int_{0}^{\boldsymbol{\cdot}}\Lambda(t)[\mathrm{d}X(t),Y(t)]-\int_{0}^{\boldsymbol{\cdot}}\Lambda(t)[X(t),\mathrm{d}Y(t)].\end{split}$ (5.1)
@@
$\mathbb{L}^{r}\text{-}\lim_{\Pi^{*}\in\mathcal{P}_{\mathbb{R}_{+}}^{*}}\sum_{t\in\Pi}\Lambda(t_{*})[X(t\wedge\cdot)-X(t_{-}\wedge\cdot),Y(t\wedge\cdot)-Y(t_{-}\wedge\cdot)]=0.$
In particular, $\mathbb{L}^{r}\text{-}\lim_{\Pi\in\mathcal{P}_{\mathbb{R}_{+}}}\mathrm{RS}_{\Pi}^{X,Y}(\Lambda)=0$.
Let $t\geq 0$, and define $C_{t}\coloneqq\sup\{\lVert\Lambda(s)\rVert_{p,q;r}:0\leq s\leq t\}<\infty$. If $\Pi^{*}\in\mathcal{P}_{\mathbb{R}_{+}}^{*}$, then
@@ 5.2 Construction of quadratic covariation
    $\Bigg\|\sum_{s\in\pi}\Lambda(s_{*})[\Delta_{s}M,\Delta_{s}N]\Bigg\|_{1}\leq\int_{0}^{t}\lVert\Lambda\rVert_{2,2;1}^{\pi^{*}}\,\mathrm{d}\kappa_{M,N}$ for all $\pi^{*}\in\mathcal{P}_{[0,t]}^{*}$.
@@
$=\sum_{s\in\pi}\lVert\Lambda(s_{*})\rVert_{2,2;1}\kappa_{M,N}((s_{-},s])=\int_{0}^{t}\lVert\Lambda\rVert_{2,2;1}^{\pi^{*}}\,\mathrm{d}\kappa_{M,N},$
@@
If $|\pi|\to 0$, then $s_{+}\searrow s$. Thus, by the continuity of $X$ and $Y$, $\Lambda(s)[\Delta_{s_{+}}X,\Delta_{s_{+}}Y]\to 0$ in $L^{1}(\mathbb{E}_{{\mathcal{C}}})$ as $|\pi|\to 0$. It therefore follows, by taking $|\pi|\to 0$, from Theorem 5.7 and Definition 5.8 that
@@ 5.3 Tools to calculate quadratic covariations
For the third claim, note that if $\Lambda$ is $\lVert\cdot\rVert_{2,2;1}$-continuous, $t\geq 0$, and $\Pi^{*}\in\mathcal{P}_{\mathbb{R}_{+}}^{*}$, then
$\sup_{0\leq s\leq t}\delta_{\Pi^{*}}(s)$ $\coloneqq\sup_{0\leq s\leq t}\Bigg\|\sum_{r\in\Pi}\Lambda(r_{*})[\Delta_{r}M^{s},\Delta_{r}N^{s}]-\mathrm{RS}_{\Pi}^{M,N}(\Lambda)(s)\Bigg\|_{1}$
@@
$=\kappa_{M,N}((0,t])\sup_{u,v\leq t:|u-v|\leq|\Pi|}\lVert\Lambda(u)-\Lambda(v)\rVert_{2,2;1}\xrightarrow[\Pi^{*}\in\mathcal{P}_{\mathbb{R}_{+}}^{*}]{|\Pi|\to 0}0$
@@ 5.4 Application: Burkholder–Davis–Gundy inequalities
Furthermore,
$\lVert x\rVert_{\mathcal{H}^{2}(\mathcal{M})}^{2}=\tau\bigg[x_{0}^{*}x_{0}+\sum_{n=1}^{N}(x_{n}-x_{n-1})^{*}(x_{n}-x_{n-1})\bigg]=\tau\bigg[x_{0}x_{0}^{*}+\sum_{n=1}^{N}(x_{n}-x_{n-1})(x_{n}-x_{n-1})^{*}\bigg].$
@@
(ii) First, observe $a\in L^{q}(\mathbb{E})$ and $\lVert a\rVert_{q}\leq\liminf_{n\to\infty}\lVert a_{n}\rVert_{q}\leq\sup\{\lVert a_{n}\rVert_{q}:n\in\mathbb{N}\}<\infty$ by the first item. For the second claim, the case $p=1$ is obvious, so we assume $p>1$. If $b\in\mathcal{A}$ and $M>0$, then
@@
Furthermore, $\lVert M(t)\rVert_{2}^{2}=\mathbb{E}\big[M^{*}(0)M(0)+[M^{*},M]_{t}^{P}\big]=\mathbb{E}\big[M(0)M^{*}(0)+[M,M^{*}]_{t}^{P}\big]$.
@@ 5.5 Examples
$\begin{split}\mu_{M,N}((u,v])\,&\mathbb{E}[\Xi[M(t)-M(s),N(t)-N(s)]\mid\mathcal{A}_{s}]\\ &=\mu_{M,N}((s,t])\,\mathbb{E}[\Xi[M(v)-M(u),N(v)-N(u)]\mid\mathcal{A}_{u}]\,\text{ and}\end{split}$ (5.15)
@@
$\begin{split}\mathbb{E}[\Xi[M(t)-M&(s),N(t)-N(s)]\mid\mathcal{A}_{s}]\\ &=\frac{\mathbb{E}[\Xi[M(s+r)-M(s),N(s+r)-N(s)]\mid\mathcal{A}_{s}]}{\mu_{M,N}((s,s+r])}\,\mu_{M,N}((s,t])\quad(s<t,\;\Xi\in\mathcal{T}_{2,s}).\end{split}$ (5.17)
@@
Above, we have considered $\mathbb{R}_{+}\ni t\mapsto X(t)\coloneqq X(t,\cdot)\in L^{2}(\tau_{n})$ as a noncommutative $L^{2}$-martingale in the usual way. Now, by the “magic formulas” ([28, §3.1]),
@@ 6.1 Adapted $C^{k}$ maps
    if $i=0,\ldots,k$, $t\geq 0$, and $(a,b)\in(\mathcal{A}_{t}\times\mathcal{B}_{t})\cap U$, then $D_{1}^{i}F(a,b)$ belongs to $\mathcal{F}_{i,t}(\mathbb{E}_{{\mathcal{A}}},\ldots,\mathbb{E}_{{\mathcal{A}}};\mathbb{E}_{{\mathcal{C}}})$; and
@@
    if $j=0,\ldots,\ell$, $t\geq 0$, and $(a,b)\in(\mathcal{A}_{t}\times\mathcal{B}_{t})\cap U$, then $D_{2}^{j}F(a,b)$ belongs to $\mathcal{F}_{j,t}(\mathbb{E}_{{\mathcal{B}}},\ldots,\mathbb{E}_{{\mathcal{B}}};\mathbb{E}_{{\mathcal{C}}})$.
@@
where $S_{k}$ is the symmetric group on $k$ letters. Thus, by Proposition 3.9, $F\in C_{a}^{\infty}(\mathcal{U};\mathcal{A})$.
@@
$\big(\partial_{\mathbf{b}_{k}}\cdots\partial_{\mathbf{b}_{1}}f_{*}(\mathbf{a}))(\omega)=\partial_{\mathbf{b}_{k}(\omega)}\cdots\partial_{\mathbf{b}_{1}(\omega)}f(\mathbf{a}(\omega))\qquad\big(\mathbf{a},\mathbf{b}_{i}\in\mathcal{A}_{n,\beta}^{d},\text{ a.e.\ }\omega\in\Omega\big).$ (6.2)
@@ 6.3 Examples: Trace smooth maps
The closure of $\{P_{{(\mathcal{A},\mathbb{E})}}:P\in(\mathrm{TrP}_{n,k,d}^{*})^{m}\}$ in $BC_{\operatorname{loc}}(\mathcal{A}_{\beta}^{n};\mathbb{B}_{k}(\mathcal{A}_{\gamma}^{d};\mathcal{A}^{m}))$ is contained in $C_{\mathbb{E}}(\mathcal{A}_{\beta}^{n};\mathbb{B}_{k}(\mathcal{A}_{\gamma}^{d};\mathcal{A}^{m}))$. In particular, if $F\in C^{k}(\mathcal{A}_{\beta}^{n};\mathcal{A}^{m})$ is such that for all $i=0,\ldots,k$, the $i^{\text{th}}$ derivative $D^{i}F$ belongs to the closure of $\{P_{{(\mathcal{A},\mathbb{E})}}:P\in(\mathrm{TrP}_{n,i,(n,\ldots,n)}^{*})^{m}\}$ in $BC_{\operatorname{loc}}(\mathcal{A}_{\beta}^{n};\mathbb{B}_{i}((\mathcal{A}_{\beta}^{n})^{i};\mathcal{A}^{m}))$, then $F\in C_{\mathbb{E}}^{k}(\mathcal{A}_{\beta}^{n};\mathcal{A}^{m})$. Consequently, the tracial noncommutative $C^{k}$ functions introduced and studied in [37] provide examples of elements of $C_{\mathbb{E}}^{k}(\mathcal{A}_{\operatorname{sa}}^{n};\mathcal{A}^{m})$. In the next section, we give examples of this kind that arise from the functional calculus (i.e., operator functions); see Remark 6.40.
@@ 6.4 Examples: Scalar functions
If $m=k+1$, $\mathbf{a}=(a_{1},\ldots,a_{k+1})\in\mathcal{M}_{\operatorname{sa}}^{k+1}$, and $\Omega_{j}=\sigma(a_{j})$ ($j=1,\ldots,k+1$) as well, then
@@
By an elementary induction argument,
$f^{[k]}(\lambda_{1},\ldots,\lambda_{k+1})=\sum_{i=1}^{k+1}f(\lambda_{i})\prod_{j\neq i}(\lambda_{i}-\lambda_{j})^{-1}$
for all distinct $\lambda_{1},\ldots,\lambda_{k+1}\in S$. In particular, $f^{[k]}$ is symmetric in its arguments. Now, we state a useful expression for $f^{[k]}$ when $f\in C^{k}(\mathbb{R})$ or when $f\colon\mathbb{C}\to\mathbb{C}$ is entire; see [49, Prop. 2.1.3(ii)] for a proof.
@@ Appendix A $L^{p}$ spaces of $\mathrm{C}^{*}$-probability spaces
Let $p\in[1,\infty)$, and write $\lVert a\rVert_{p}\coloneqq\mathbb{E}[|a|^{p}]^{1/p}$ and $\lVert a\rVert_{\infty}\coloneqq\lVert a\rVert$ for $a\in\mathcal{A}$.
@@
    If $a\in\mathcal{A}$, then $\lVert a\rVert_{p}=\sup\{|\mathbb{E}[ab]|:b\in\mathcal{A},\,\lVert a\rVert_{q}\leq 1\}$. If $1<p,q<\infty$ as well, then the map $\mathcal{A}\ni a\mapsto(b\mapsto\tilde{\mathbb{E}}[ab])\in L^{p}(\mathcal{A},\mathbb{E})^{*}$ extends to an isometric isomorphism $L^{q}(\mathcal{A},\mathbb{E})\to L^{p}(\mathcal{A},\mathbb{E})^{*}$.
@@ Appendix B Notation index
$\ell_{(\operatorname{loc})}^{\infty}(\mathbb{R}_{+};\mathcal{V})$ (locally) bounded functions $\mathbb{R}_{+}\to\mathcal{V}$; pp. 4.5 & 5.7